OpenStudy (anonymous):

How many milligrams of C atoms are in 2.04 x 10^20 molecules of ethanol?

OpenStudy (zepp):

Ethanol: \(\large \text{C}_2\text{H}_6\text{O}\)

OpenStudy (zepp):

So first question, how many moles does 2.04 x 10^20 molecules represent?

OpenStudy (anonymous):

its not CH(Sub 6)O?

OpenStudy (anonymous):

why did u give Carbon2?

OpenStudy (zepp):

Because that's the formula for Ethanol

OpenStudy (zepp):

It's a 2-carbon alcohol.

OpenStudy (anonymous):

\[(6.02 \times 10^{23})(2.04 \times 10^{20})?\]

OpenStudy (anonymous):

oh ok

OpenStudy (zepp):

That's incorrect; \[\large 1 \text{ mole}= 6.02×10^{23}\]

OpenStudy (zepp):

\[\large ? \text{ mole}=2.04×10^{20}\]

OpenStudy (zepp):

That's a proportion and you should have this: \(\LARGE \frac{2.04×10^{20}}{6.02×10^{23}}\)

OpenStudy (anonymous):

oh so its divide instead of multiply?

OpenStudy (zepp):

Yes, that's algebraic manipulation.

OpenStudy (zepp):

Once you get that, you'll get the number of moles this quantity of molecules represent, are afterward, you may convert that easily to its mass through the molar mass.

OpenStudy (anonymous):



OpenStudy (zepp):

Why 1000?

OpenStudy (anonymous):

for milligrams?

OpenStudy (zepp):

Our units are in moles...

OpenStudy (zepp):

\[\large \frac{1 \text{ mole}}{6.02*10^{23} \text{ atoms}}=\frac{x \text{ mole}}{2.04 * 10^{20}\text{ atoms}}\]Multiply both sides by 2.04*10^20\[\large \frac{2.04*10^{20}\text{ atoms}*1 \text{ mole}}{6.02*10^{23} \text{ atoms}}=x \text{ mole}\]

OpenStudy (zepp):

atoms and atoms simplify out, leaving the mole on both sides.\[\large \frac{2.04*10^{20}}{6.02*10^{23}}\text{mole}=x \text{ mole}\]

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

i got .0003 correct?

OpenStudy (zepp):

You'll get as answer: \(\large x\text{ moles}\) Now you may proceed to transform this using \(\LARGE n=\frac{m}{M}\), where \(\large \text{n}\) is the number of moles, \(\large \text{m}\) is the mass in \(\large \textbf{grams}\) (so you'll need to retransform it into milligrams as you're asked to do) and \(\large \text{M}\) is the molar mass.

OpenStudy (anonymous):

Molar mass is 46.07?

OpenStudy (zepp):

Yes

OpenStudy (anonymous):

well for ethanol it self right?

OpenStudy (anonymous):

would i multiply that by .0003

OpenStudy (zepp):

Exact.

OpenStudy (anonymous):

46.68->46.7 sorry

OpenStudy (anonymous):

so i got .0140 because i used 46.68*.0003

OpenStudy (anonymous):

thats just grams right?

OpenStudy (zepp):

Yep

OpenStudy (anonymous):

.0140 x 10^-3 ?

OpenStudy (zepp):

Good, now you'll need to set up another proportion:\[ \frac{\text{Mass of Carbon in 1 mole of Ethanol}}{\text{Molar mass of Ethanol}}=\frac{\text{Mass of Carbon (Unknown) }}{\text{Mass of the Ethanol given to you}}\]

OpenStudy (anonymous):

(?/46.68)=(?/.0140x10^-3?)

OpenStudy (zepp):

Mass of carbon in 1 mole of ethanol is a known number, 2*Molar mass of carbon, 12, 2*12 = 24.

OpenStudy (anonymous):

oh ok i see where you got that from

OpenStudy (anonymous):

(24/46.68) = (?/.0140x10^-3?)

OpenStudy (anonymous):

was i correct on the second half too?

OpenStudy (zepp):

I don't know, I didn't push the values in my calculator, I've shown you how to do these kind of problem, all the other mathematical manipulations/calculations should be done by you; the right method is shown so you can take down any question that looks like this.

OpenStudy (anonymous):

I was asking about for the mass of the ethanol that the answer i found before right?

OpenStudy (zepp):

Like I said, I don't know :)

OpenStudy (anonymous):

Ok so now i simple mult across top and bottom and cancel out to get my answer

OpenStudy (anonymous):

simply*

OpenStudy (anonymous):

i think i have it

OpenStudy (anonymous):

thank you

OpenStudy (zepp):

np